OpenStudy (anonymous):

14. 4^7 over 4^9 (1 point) –16 one sixteenth one over eight 8

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm 4^7\div4^9=4^{7-9}=4^{-2}=1/4^2=1/16 }\)